# Multiplicity of young isolated planetary mass objects in Taurus and Upper Scorpius

### v1

@@ Abstract
Methods. We use the Hubble Space Telescope and its Wide Field Camera 3 and the Very Large Telescope and its ERIS adaptive optics facility to search for visual companions among a sample of 77 objects members of the Upper Scorpius and Taurus young nearby associations with estimated masses in the range between approximately 5–33 MJup .
Results. We report the discovery of one companion candidate around a Taurus member with a separation of 111.90.4 mas, or 18 au assuming a distance of 160 pc, with an estimated primary mass in the range between 3–6 MJup and a secondary mass between 2.6–5.2 MJup , depending on the assumed age. This corresponds to an overall binary fraction of 1.8% among free-floating planetary mass objects over the separation range 7 au. Despite the limitations of small-number statistics and variations in spatial resolution and sensitivity, our results, combined with previous high-spatial-resolution surveys, suggest a notable difference in the multiplicity properties of objects below 25 MJup between Upper Sco and Taurus. In Taurus, five companions were identified among 78 observed objects (4.9%), whereas none were found among 97 objects in Upper Sco (1.2%).
@@
Free-floating planetary mass objects (FFPs), also called isolated planetary mass-objects (IPMOs), are planetary-mass objects that do not orbit a star, but roam the galaxy isolated. They are some of the most challenging astrophysical objects to study: incapable of sustaining nuclear fusion, they are intrinsically extremely faint and steadily fade over time to become very difficult to detect directly. Young FFPs with super-Jovian masses in star forming regions had been discovered using early direct imaging (e.g. Tamura et al., 1998; Lucas & Roche, 2000; Zapatero Osorio et al., 2000). In contrast, less-massive FFPs had been discovered using indirect methods: gravitational micro-lensing surveys have proven particularly successful to detect them down to a few Earth masses (e.g. Mróz et al., 2017, 2020; Ryu et al., 2021; Sumi et al., 2023, and references therein). The future Roman Space Telescope will be sensitive to FFP down to masses as low as 0.1 M⊕ and is predicted to harvest tens of such objects during the course of its micro-lensing survey (Johnson et al., 2020). The advent of large and very deep ground-based surveys led to the discovery of several hundreds FFP candidates in nearby star forming regions (Peña Ramírez et al., 2012; Scholz et al., 2012; Mužić et al., 2011, 2015; Bayo et al., 2011; Esplin & Luhman, 2019; Miret-Roig et al., 2022, and references therein) and the field (Delorme et al., 2008, 2017). The WISE survey (Wright et al., 2010) and its subsequent extensions led to the discovery of several FFPs in the solar neighborhood (Beichman et al., 2013; Schneider et al., 2016). More recently the JWST and Euclid missions are starting to reveal large samples of FFPs thanks to their unprecedented sensitivities in the infrared (Bouy, H. et al., 2025; Luhman et al., 2024; Martín et al., 2024; Langeveld et al., 2024; De Furio et al., 2024, 2025).
The origin and the formation of FFP remains poorly understood. Even though they represent a very small fraction of the total baryonic mass of our galaxy, they add up to at least 2 to 5% of the astrophysical objects populating the Milky Way (Miret-Roig et al., 2022), making them more numerous than stars with masses greater than 3 M⊙. Similar to slightly more massive brown dwarfs, several scenarios are considered to explain their existence:
@@
While recent direct observational evidence confirms that these different processes are at work (Miret-Roig et al., 2022; Palau et al., 2024; Bouy et al., 2009, and references therein), we still do not understand their relative contributions to the overall FFP population (e.g., photo-erosion can only occur in the direct vicinity of relatively scarce OB stars) and how the environment affects their outcome.
@@ 2.1 HST targets
We selected 30 targets in the USco sample of FFPs from Miret-Roig et al. (2022), and 30 targets in the sample of Taurus brown dwarfs and FFP candidates from Esplin & Luhman (2019) and Bouy et al. (in prep.). The latter used the same methodology and strategy as Miret-Roig et al. (2022) based on proper motions and multi-wavelength photometry in the Taurus molecular clouds to identify high-probability candidate members down to a few Jupiter masses only. Figure 1 shows (i, i-z) colour-magnitude diagrams of the targets in each region. In USco, we randomly selected targets below the planetary mass limit plus 3 mag of extinction, to minimize contamination by reddened more massive objects. The same strategy was initially used for the Taurus sample, but the lack of suitable HST guide stars in the much sparser Taurus region forced us to drop some targets in favor of slightly more massive objects with guide stars available.
Recent Gaia results have revealed that the various subgroups within these two associations span a wide range of distances–from 130 to 200 pc, with an average around 140 pc–and cover ages of approximately 1–10 Myr in Taurus and 3–20 Myr in USco (Galli et al., 2019; Ratzenböck et al., 2023). Our targets lack sufficient data (specifically parallax and radial velocity measurements) to determine their precise subgroup membership. Therefore, for the purposes of this study and in all subsequent figures, we adopt an average distance of 140 pc for both samples, and assume ages of 3 Myr for the Taurus targets and 5 Myr for those in Upper Sco.
The USco targets thus sample the mass range between 611 MJup and the Taurus sample between 533 MJup , estimated following the method described in Miret-Roig et al. (2022) by analyzing the available multi-wavelength photometry with Sakam (Romero, 2019) and the ages and distances mentioned above. Figures 2 and 3 show that the targets are randomly distributed in each association and form a representative sample of the overall associations and the sub-groups composing them.
@@ 2.2 VLT targets
We selected 60 targets from the USco free-floating planet sample identified by Miret-Roig et al. (2022), distinct from the HST sample described above, based upon the availability of a sufficiently bright reference star for the adaptive optics wavefront sensor within 15″. The selection was also designed to cover the 12–30 MJup mass range, complementing the HST sample mentioned earlier and prior brown dwarf multiplicity studies in USco by Bouy et al. (2006); Kraus et al. (2006); Biller et al. (2011); Kraus & Hillenbrand (2012).
@@ 3.2 VLT ERIS/NIX
Observations with the Enhanced Resolution Imager and Spectrograph (ERIS) adaptive optics (AO) instrument and its NIX imager (Davies et al., 2023) were conducted at the VLT in queue mode for 25 of the 60 targets between April and December 2023 as part of program 111.24H3 (P.I. Bouy). However, due to insufficient data quality in 3 cases, the final dataset comprised 22 objects. Table 5 gives a summary of their properties as reported in Miret-Roig et al. (2022). Figure 2 shows their positions in the association, and Figure 1 shows their positions in a (i, i-z) color-magnitude diagrams. The targets were too faint for the wavefront sensor and nearby (15″) bright (Gaia RP16.5 mag) stars were used as reference for the adaptive optics. NIX has a pixel scale of 13 mas, and the observations were conducted in the J band, which provides an optimal balance between sensitivity to the cool, red targets and spatial resolution. Each target was observed with 5 DIT of 5 s exposures at 9 jittered positions, resulting in a total on-source exposure time of 225 s per target. Raw images and associated calibration frames were retrieved from the ESO archive, and the data was processed using the official ESO ERIS pipeline (v1.7.0). Table 5 reports the Strehl ratios of the observations, ranging between 2 and 32%.
@@ 4.1 HST images
We estimated component masses by linearly interpolating the predicted values from the Chabrier et al. (2023) models for ages of 1Myr and 3Myr, assuming a distance of 160pc. The resulting mass range for each component is listed in Table2. These values should be considered lower limits, as extinction could increase them. For instance, assuming a visual extinction of mag would yield a primary mass range of 5–10 MJup and a secondary mass range of 4.5–9 MJup . In any case, the mass ratio is relatively high, approximately 0.9.
@@ 4.1.2 StraKLIP
Candidate validation: for non visual binaries, jack-knife tests are perform to insure that these candidates are not artifacts introduced by one of the reference images. This test is performed by subtracting the target with a PSF library where one at the time one reference is removed, and by analyzing each single residual. If the candidate disappears in one of them, that’s a clear sign that was injected, and therefore not real.
@@ 4.1.3 PUBLIC-WiFi
PUBLIC-WiFi (Previously Undiscovered Binaries Lying In Clusters - Wide Field) 222https://github.com/aggle/public_wifi, like StraKLIP, is a principal component analysis (PCA)-based faint companion detection pipeline designed to organize the analysis of large wide-field surveys of unocculted point sources. The main points of difference are in architecture – PUBLIC-WiFi is primarily object-oriented where StraKLIP is primarily database-oriented – and in some specifics of implementation, e.g. how PSF forward modeling is implemented (Wang et al., 2015a; Pueyo, 2016; Ruffio et al., 2017). Though the two algorithms necessarily share many of the same basic steps described in 4.1.2, they nevertheless provide semi-independent checks on each other with regard to companion detection and search depth.
@@ 4.2 VLT images
Instead, we limited our analysis to a visual inspection of the images. Figure 8 shows that, in most cases, our observations should have also been sensitive enough to detect companions with a flux ratio of 0.5 (J=0.75 mag) at separations as small as 005 (or approximately 7 au at the distance of Taurus and USco). This sensitivity is comparable to that achieved in previous studies of the same association for similar ultracool objects by Bouy et al. (2006); Kraus et al. (2006); Biller et al. (2011); Kraus & Hillenbrand (2012) and to the HST images presented above. This inspection did not reveal any companions within 06 of the 22 targets. Companions at larger separations would have been detected and reported in Miret-Roig et al. (2022) as common proper-motion companions.
@@ 5.2 Dependence on environmental conditions
In that context, Figure 10 presents (i, i-z) color-magnitude diagrams for objects observed at high spatial resolution in Taurus and USco using HST or adaptive optics over the past two decades, as reported in the present work and Todorov et al. (2010, 2014); Bouy et al. (2006); Biller et al. (2011); Kraus & Hillenbrand (2012); Kraus et al. (2006); Konopacky et al. (2007); Luhman et al. (2007, 2009). Both regions have been extensively studied, with statistically significant samples including over 100 sources between a few Jupiter masses and 50 MJup in each association. The spatial resolution and sensitivity of these HST and AO studies are variable but comparable, typically achieving flux ratios of 0.5 or better at separations of about 005008.
Since Taurus and USco are located at similar distances (approximately 140pc) and have comparable ages, the sensitivity of all these surveys to BD and FFP and their companions is effectively very similar for both regions. Notably, no companions have been detected in USco among the 98 objects with masses below 25MJup , corresponding to an upper limit on the binary fraction of up to 1.2%33footnotemark: 3. In contrast, so far four binaries have been identified in Taurus among the 78 objects observed at high spatial resolution within the same mass range, corresponding to a binary fraction of 4.9% and highlighting a potentially significant difference in the population of low-mass companions between the two regions. Above 25 MJup the relative numbers of multiple systems identified are statistically indistinguishable between the two associations, as already reported in the previous studies mentioned above (Bouy et al., 2006; Biller et al., 2011; Kraus & Hillenbrand, 2012; Kraus et al., 2006; Konopacky et al., 2007).
5.3 Implications for the formation process of FFPs
@@
The Taurus molecular clouds and USco were chosen for this study because they represent drastically different environments (Reipurth, 2008a, b). While USco includes several massive B stars and around 3 500 members (Blaauw, 1946; Miret-Roig et al., 2022), Taurus is a sparser T-association with no massive star and around 900 low-mass members located around/within the filamentary clouds (Galli et al., 2019; Joncour et al., 2018). USco is also slightly older, with groups ranging from 3 to 19 Myr compared to 1 to 10 Myr for Taurus.
@@
If real, the apparent lack of binaries below 25 MJup in USco could be the result of different formation pathways. A variety of formation channels for binary FFPs have been proposed, including both dynamic ejection mechanisms such as planet-planet (or BD-BD) scattering Veras & Raymond (2012); Lazzoni et al. (2024), stellar flybys Wang et al. (2024), as well as other mechanisms such as photo-erosion of fragmenting stellar cores Diamond & Parker (2024) or via grazing encounters between protoplanetary disks Fu et al. (2025). Very dynamic formation channels likely produce a lower binary population than calmer ones, and it is possible that the dominant formation channels in USco and Taurus are different. Conversely, it could be that fewer objects in the 25MJup mass range form through core-collapse in USco compared to Taurus. Core-collapse is indeed generally thought to be the main formation mechanism for these binaries, and it may be challenging to form very low-mass cores that could eventually become FFPs or low-mass BDs in denser regions with strong feedback from massive stars and their supernovæ (Preibisch & Zinnecker, 1999).
The difference between the two regions could instead be the result of dynamical disruption of binary FFPs by stellar encounters Parker & Quanz (2012); Yu & Lai (2024); Wang et al. (2015b); Huang et al. (2024). Simulations by Portegies Zwart & Hochart (2024) show that binary FFPs in an USco-like environment are disrupted with a half life Myr. Unless they are continuously resupplied, it would not be surprising that the older and denser cluster would have a lower binary rate.

### v2

# Multiplicity of young brown dwarfs and isolated planetary mass objects in Taurus and Upper Scorpius

@@ Abstract
Methods. We use the Hubble Space Telescope and its Wide Field Camera 3 and the Very Large Telescope and its ERIS adaptive optics facility to search for visual companions among a sample of 77 objects members of the Upper Scorpius and Taurus young nearby associations with estimated masses in the range between approximately 6–66 MJup .
Results. We report the discovery of one companion candidate around a Taurus member with a separation of 111.90.4 mas, or 18 au assuming a distance of 160 pc, with an estimated primary mass in the range between 3–6 MJup and a secondary mass between 2.6–5.2 MJup , depending on the assumed age. This corresponds to an overall binary fraction of 1.8% among low-mass brown dwarfs and free-floating planetary mass objects over the separation range 7 au. Despite the limitations of small-number statistics and variations in spatial resolution and sensitivity, our results, combined with previous high-spatial-resolution surveys, suggest a notable difference in the multiplicity properties of objects below 30–50 MJup between Upper Sco and Taurus. In Taurus, a binary fraction of % is found for objects with masses below 30MJup , and of % for objects with masses below 50MJup , whereas no binary were found among 80 objects over the matching luminosity range in Upper Sco, corresponding to an upper limit of 1.2%.
Conclusions. This difference may point to intrinsically distinct formation conditions, with warmer parental molecular clouds originally present in Upper Sco potentially inhibiting fragmentation into the lowest-mass brown dwarfs and free-floating planets compared to cooler environments such as Taurus.
@@
Free-floating planetary mass objects (FFPs), also called isolated planetary mass-objects (IPMOs), are planetary-mass objects that do not orbit a star, but roam the galaxy isolated. They are some of the most challenging astrophysical objects to study: incapable of sustaining nuclear fusion, they are intrinsically extremely faint and steadily fade over time to become very difficult to detect directly. Young FFPs with super-Jovian masses in star forming regions had been discovered using early direct imaging (e.g. Tamura et al., 1998; Lucas & Roche, 2000; Zapatero Osorio et al., 2000). In contrast, less-massive FFPs had been discovered using indirect methods: gravitational micro-lensing surveys have proven particularly successful to detect them down to a few Earth masses (e.g. Mróz et al., 2017, 2020; Ryu et al., 2021; Sumi et al., 2023, and references therein). The future Roman Space Telescope will be sensitive to FFP down to masses as low as 0.1 M⊕ and is predicted to harvest tens of such objects during the course of its micro-lensing survey (Johnson et al., 2020). The advent of large and very deep ground-based surveys led to the discovery of several hundreds FFP candidates in nearby star forming regions (Peña Ramírez et al., 2012; Scholz et al., 2012; Mužić et al., 2011, 2015; Bayo et al., 2011; Esplin & Luhman, 2019; Miret-Roig et al., 2022a, and references therein) and the field (Delorme et al., 2008, 2017). The WISE survey (Wright et al., 2010) and its subsequent extensions led to the discovery of several FFPs in the solar neighborhood (Beichman et al., 2013; Schneider et al., 2016). More recently the JWST and Euclid missions are starting to reveal large samples of FFPs thanks to their unprecedented sensitivities in the infrared (Bouy, H. et al., 2025; Luhman et al., 2024; Martín et al., 2024; Langeveld et al., 2024; De Furio et al., 2024, 2025b).
The origin and the formation of FFP remains poorly understood. Even though they represent a very small fraction of the total baryonic mass of our galaxy, they add up to at least 2 to 5% of the astrophysical objects populating the Milky Way (Miret-Roig et al., 2022a), making them more numerous than stars with masses greater than 3 M⊙. Similar to slightly more massive brown dwarfs, several scenarios are considered to explain their existence:
@@
While recent direct observational evidence confirms that these different processes are at work (Miret-Roig et al., 2022a; Palau et al., 2024; Bouy et al., 2009, and references therein), we still do not understand their relative contributions to the overall FFP population (e.g., photo-erosion can only occur in the direct vicinity of relatively scarce OB stars) and how the environment affects their outcome.
@@ 2.1 HST targets
We selected 30 targets in the USco sample of FFPs from Miret-Roig et al. (2022a), and 30 targets in the sample of Taurus brown dwarfs and FFP candidates from Esplin & Luhman (2019) and Bouy et al. (in prep.). The latter used the same methodology and strategy as Miret-Roig et al. (2022a) based on proper motions and multi-wavelength photometry in the Taurus molecular clouds to identify high-probability candidate members down to a few Jupiter masses only. Figure 1 shows (i, i-z) colour-magnitude diagrams of the targets in each region. Because both samples are based on a selection using both proper motions and multi-wavelength photometry and given the relatively large mean proper motions of Taurus and USco members with respect to field objects, the contamination rate by background or foreground sources unrelated to theses star forming regions is expected to be very low (Bouy et al., 2022).
In USco, we randomly selected targets below the planetary mass limit plus 3 mag of extinction, to minimize contamination by reddened more massive objects. The same strategy was initially used for the Taurus sample, but the lack of suitable HST guide stars in the much sparser Taurus region forced us to drop some targets in favor of slightly more massive objects with guide stars available.
Recent Gaia results have revealed that the various subgroups within these two associations span a wide range of distances (from 130 to 200 pc, with an average around 140 pc) and ages, but overlap on the plane of the sky and in the proper motion space. The Taurus association includes groups with ages ranging mostly from 1 to 3Myr (Luhman et al., 2023), while USco includes sub-groups with ages between 3–20 Myr, with a median age around 10 Myr (Schmitt et al., 2022; Miret-Roig et al., 2022b; Luhman, 2022; Žerjal et al., 2023; Ratzenböck et al., 2023). Our targets lack sufficient data (specifically parallax and radial velocity measurements) to determine their precise subgroup membership. Therefore, for the purposes of this study and in all subsequent figures, we adopt an average distance of 140 pc for both samples, assume an age of 3 Myr for the Taurus targets and 5–10 Myr for those in Upper Sco.
The USco targets thus sample the mass range between 6–66 MJup and the Taurus sample between 5–33 MJup , estimated following the method described in Miret-Roig et al. (2022a) by analyzing the available multi-wavelength photometry with Sakam (Romero, 2019) and the ages and distances mentioned above. Figures 2 and 3 show that the targets are randomly distributed in each association and form a representative sample of the overall associations and the sub-groups composing them.
@@ 2.2 VLT targets
We selected 60 targets from the USco free-floating planet sample identified by Miret-Roig et al. (2022a), distinct from the HST sample described above, based upon the availability of a sufficiently bright reference star for the adaptive optics wavefront sensor within 15″. The selection was also designed to cover the 12–30 MJup mass range, complementing the HST sample mentioned earlier and prior brown dwarf multiplicity studies in USco by Bouy et al. (2006); Kraus et al. (2006); Biller et al. (2011); Kraus & Hillenbrand (2012).
@@ 3.2 VLT ERIS/NIX
Observations with the Enhanced Resolution Imager and Spectrograph (ERIS) adaptive optics (AO) instrument and its NIX imager (Davies et al., 2023) were conducted at the VLT in queue mode for 25 of the 60 targets between April and December 2023 as part of program 111.24H3 (P.I. Bouy). However, due to insufficient data quality in 3 cases, the final dataset comprised 22 objects. Table 5 gives a summary of their properties as reported in Miret-Roig et al. (2022a). Figure 2 shows their positions in the association, and Figure 1 shows their positions in a (i, i-z) color-magnitude diagrams. The targets were too faint for the wavefront sensor and nearby (15″) bright (Gaia RP16.5 mag) stars were used as reference for the adaptive optics. NIX has a pixel scale of 13 mas, and the observations were conducted in the J band, which provides an optimal balance between sensitivity to the cool, red targets and spatial resolution. Each target was observed with 5 DIT of 5 s exposures at 9 jittered positions, resulting in a total on-source exposure time of 225 s per target. Raw images and associated calibration frames were retrieved from the ESO archive, and the data was processed using the official ESO ERIS pipeline (v1.7.0). Table 5 reports the Strehl ratios of the observations, ranging between 2 and 32%.
@@ 4.1 HST images
We estimated component masses by linearly interpolating the predicted values from the Chabrier et al. (2023) models for ages of 1Myr and 3Myr, assuming a distance of 160pc. The resulting mass range for each component is listed in Table2. These values should be considered lower limits, as extinction could increase them. The 3D extinction map of Green et al. (2019) reports an integrated line-of-sight extinction of 1.1 mag at the position of J042705.86+261520.3. Assuming a conservative visual extinction of mag would yield a primary mass range of 5–10 MJup and a secondary mass range of 4.5–9 MJup . In any case, the mass ratio is relatively high, approximately 0.9.
@@ 4.1.2 StraKLIP
Candidate validation: for non visual binaries, jack-knife tests are performed to insure that these candidates are not artifacts introduced by one of the reference images. This test is performed by subtracting the target with a PSF library where one at the time one reference is removed, and by analyzing each single residual. If the candidate disappears in one of them, that’s a clear sign that was injected, and therefore not real.
@@ 4.1.3 PUBLIC-WiFi
PUBLIC-WiFi (Previously Undiscovered Binaries Lying In Clusters - Wide Field)222https://github.com/aggle/public_wifi, like StraKLIP, is a principal component analysis (PCA)-based faint companion detection pipeline designed to organize the analysis of large wide-field surveys of unocculted point sources. The main points of difference are in architecture – PUBLIC-WiFi is primarily object-oriented where StraKLIP is primarily database-oriented – and in some specifics of implementation, e.g. how PSF forward modeling is implemented (Wang et al., 2015; Pueyo, 2016; Ruffio et al., 2017). Though the two algorithms necessarily share many of the same basic steps described in 4.1.2, they nevertheless provide semi-independent checks on each other with regard to companion detection and search depth.
@@ 4.2 VLT images
Instead, we limited our analysis to a visual inspection of the images. Figure 8 shows that, in most cases, our observations should have also been sensitive enough to detect companions with a flux ratio of 0.5 (J=0.75 mag) at separations as small as 005 (or approximately 7 au at the distance of Taurus and USco). This sensitivity is comparable to that achieved in previous studies of the same association for similar ultracool objects by Bouy et al. (2006); Kraus et al. (2006); Biller et al. (2011); Kraus & Hillenbrand (2012) and to the HST images presented above. This inspection did not reveal any companions within 06 of the 22 targets. Companions at larger separations would have been detected and reported in Miret-Roig et al. (2022a) as common proper-motion companions.
@@ 5.2 Dependence on environmental conditions
In this context, we complemented the present work by compiling a comprehensive list of all low-mass sources in Taurus and Upper Sco that have been observed at high spatial resolution with either HST or AO over the past two decades. Both regions have been extensively surveyed, yielding statistically significant samples (100 sources each) that reach down to a few Jupiter masses in each association. Tables 6 and 7 list the Taurus and Upper Sco members reported in the literature in Todorov et al. (2010, 2014); Bouy et al. (2006); Biller et al. (2011); Kraus & Hillenbrand (2012); Kraus et al. (2006); Konopacky et al. (2007); Luhman et al. (2007) and Luhman et al. (2009). The corresponding colour–magnitude diagrams are shown in Fig. 10.
Some sources appear in multiple studies, occasionally under different designations. These duplications were identified and accounted for when compiling the final samples presented in Tables 6 and 7.
The spatial resolution and sensitivity of these HST and AO studies are variable but comparable, typically achieving flux ratios of 0.5 or better at separations of about 005008. Since Taurus and Upper Sco lie at comparable distances (140 pc), the high–spatial-resolution surveys performed in both regions reach very similar sensitivities to BDs, FFPs, and their companions. The interpretation of non-detections in USco, however, depends on the assumed age of that association. Given the uncertainty and spread on the ages of the USco sub-groups mentioned above, we consider two commonly adopted age scenarios: 5 Myr and 10 Myr bracketing most of the sub-groups present in that region.
(i) Case 1: If USco is assumed to be 5 Myr old, the faintest magnitude at which no companions have been detected ( mag) corresponds to a mass limit of 29 MJup according to the evolutionary models of Chabrier et al. (2023). At the younger age of 3 Myr appropriate for Taurus, this same mass corresponds to a limit of mag. Among the Taurus objects surveyed at high spatial resolution, 4 binaries are found among 67 sources with mag (see Table 6), yielding a binary fraction of %. With the exception of our new candidate (which still lacks spectroscopic confirmation), these Taurus objects have reported spectral types between M5.5 and M8.5 and therefore cannot be explained as more massive embedded objects.
(ii) Case 2: If Upper Sco is instead assumed to be 10 Myr old, the non-detection limit of mag translates to a higher mass threshold of 50 MJup. At 3 Myr in Taurus, this mass corresponds to mag. In this regime, 8 binaries are identified among 94 Taurus objects fainter than this limit, corresponding to a binary fraction of %. Again, their reported spectral types (later than M5.25) are consistent with masses below approximately 50 MJup, and thus these systems are not embedded more massive interlopers.
We note that even if we conservatively exclude our newly reported candidate until spectroscopic confirmation is available, the corresponding fractions change only marginally to % (5 Myr/29 MJup regime) and % (10 Myr/50 MJup regime). Similarly, Kerr et al. (2023) found that the multiple system [MDM2001] CFHT-BD-Tau 18 (i=15.73 mag) might be a member of the 11 Myr SCYA 108 group. If we conservatively exclude it as well, the corresponding fractions change marginally only when assuming 10 Myr to % (50 MJup regime).
In contrast with the several detections in Taurus, no visual binary candidates have been found among the 80 USco members with mag observed at high spatial resolution. This implies an upper limit on the binary fraction of % in USco,33footnotemark: 3 significantly lower than in Taurus for either of the two ages (respectively mass regimes) considered. These results therefore suggest a genuine and statistically significant difference in the multiplicity properties of objects below approximately 30–50 MJup between the two regions.
We also note that the Taurus fraction mentioned above should be considered as a lower limit: while the multiple systems have observed spectral types consistent with masses below approximately 30–50 MJup, the other sources could include a few more massive embedded interlopers that might decrease the final binary fraction in the correspond mass range.
5.3 Implications for the formation process of BDs and FFPs
@@
The Taurus molecular clouds and Upper Sco were selected for this study because they represent drastically different star-forming environments (Reipurth, 2008a, b). Upper Sco is a rich OB association hosting several massive B stars and approximately 3 500 members (Blaauw, 1946; Miret-Roig et al., 2022a), whereas Taurus is a sparse association devoid of massive stars, with about 900 low-mass members distributed along filamentary molecular clouds (Galli et al., 2019; Joncour et al., 2018). Upper Sco is also moderately older, with subgroups spanning ages from 3 to 20 Myr and a median age around 10 Myr (Miret-Roig et al., 2022b; Luhman, 2022; Schmitt et al., 2022; Žerjal et al., 2023; Ratzenböck et al., 2023), compared to typical ages of 1–3 Myr for the Taurus molecular clouds (Luhman et al., 2023).
@@
If real, the apparent lack of binary systems below 30–50 MJup in Upper Sco may therefore reflect differences in formation pathways between the two regions. A variety of mechanisms have been proposed for the formation of low-mass brown dwarfs and free-floating planets, including dynamical processes such as planet–planet or brown dwarf–brown dwarf scattering (Veras & Raymond, 2012; Lazzoni et al., 2024), stellar flybys (Wang et al., 2024), photo-erosion of fragmenting stellar cores (Diamond & Parker, 2024), or grazing encounters between protoplanetary disks (Fu et al., 2025). More dynamically active formation channels are generally expected to produce fewer bound systems than quiescent collapse scenarios, and it is plausible that the dominant mechanisms operating in Upper Sco and Taurus differ. Alternatively, the efficiency of forming very low-mass objects through core collapse may itself be reduced in Upper Sco compared to Taurus. In particular, producing extremely low-mass prestellar cores may be more difficult in environments subject to stronger radiative and mechanical feedback from massive stars and their supernovae (Preibisch & Zinnecker, 1999).
A physically motivated explanation for these differences may be linked to the thermal properties of the parental molecular clouds and their impact on fragmentation. Both the characteristic mass scale of fragments and their typical separations are expected to depend sensitively on the cloud temperature through the Jeans formalism. Since the Jeans mass scales approximately as (e.g. Eq. 14 in Palau et al., 2024), implying that warmer molecular clouds fragment at systematically higher minimum masses (assuming similar densities for opacity-limited fragmentation in all clouds). As a result, the production of very low-mass objects through core collapse, including low-mass BDs and FFPs, is expected to be suppressed in regions with elevated cloud temperatures. Palau et al. (2024) argue that this effect explains the paucity of proto–brown dwarfs near the planetary-mass limit in Ophiuchus (compared to Taurus, see their Fig. 7 and Sect. 7.4), a region known to host relatively warm and dense gas (Alves et al., 2025). Given the presence of massive stars in Upper Sco, it is likely that its parental clouds were significantly warmer than those in Taurus, which could naturally account for the reduced number of companions in the planetary-mass regime.
In addition to influencing the characteristic mass scale, the cloud temperature also affects the typical spatial scale of fragmentation. The Jeans length, which represents the minimum size of a perturbation capable of gravitational collapse, scales as (Bontemps et al., 2010). Higher temperatures therefore lead to larger initial separations between fragments. For instance, at a density of cm-3, a cloud at K (typical of dense gas as found in Ophiuchus and Orion, Alves et al., 2025; Tang et al., 2018) yields a Jeans length of approximately 280 au, compared to about 140 au at K (typical of the Taurus molecular clouds, del Burgo & Laureijs, 2005). Such a factor of two to three increase in characteristic separation may be sufficient to prevent newly formed low-mass fragments from remaining gravitationally bound, thereby reducing the likelihood of forming binary systems.
Numerical simulations further illustrate the strong sensitivity of multiplicity outcomes to the initial cloud temperature. Riaz et al. (2014) showed that, for modest azimuthal density perturbations, increasing the initial temperature above 10 K can suppress binary formation altogether, leading instead to a single low-mass fragment. Conversely, small temperature variations below 10 K can result in changes of order 100 au in the separation of binary components. These results demonstrate that relatively small differences in the thermal conditions of star-forming clouds can have a decisive impact on the formation and survival of low-mass binary systems.
Taken together, these considerations suggest that the absence of binaries below 50 MJup in Upper Sco may reflect intrinsically different initial conditions compared to Taurus, rather than observational biases alone. Warmer parental clouds would both inhibit the formation of low-mass BD and FFPs via core collapse, the main mechanism to produce binaries, and favor wider initial separations, reducing the probability that such low-mass fragments remain bound. While this interpretation remains speculative, it provides a physically motivated framework that can be tested with future observational constraints on cloud temperatures and with spectroscopic confirmation of candidate systems.